IfSubStringInString “ELL” is in “hellohello” No “ELL” in “hellohello”

BeforeSubString and BehindSubString Text befor “ell” in “hellohello”: h Text after “ell” in “hellohello”: ohello hellohello

CountSubStrings Count of “ell” in “hellohello”: 2 Count of “ell” in “hellohellohello”: 3 Count of “ell” in “hellohellohellohello”: 4

SubStringsToCounter Count of “ell” in hellohello” (counter): 2